OpenStudy (anonymous):

find the vertex of the parabola by completing the square. y=x^2+4x-8

OpenStudy (alexandervonhumboldt2):

y=x^2+4x-8 y=x^2+4x+4-8-4 y=(x+2)^2-12 vertex=(-2, 12)

OpenStudy (alexandervonhumboldt2):

woops sorry vertex = (-2, -12)

OpenStudy (anonymous):

Thanks and thank you so much for explaining how to solve it

OpenStudy (alexandervonhumboldt2):

np

OpenStudy (alexandervonhumboldt2):

may i have a medal and fan/ thank you